OpenStudy (anonymous):

The legislature of Puerto Rico consists of a 27-member Senate and a 51-member House of Representatives. How many ways are there to choose a group of seven members from the Puerto Rican legislature?

OpenStudy (anonymous):

@Hunus

OpenStudy (anonymous):

@satellite73

OpenStudy (loser66):

hello honey, your question is ambiguous since there is no rule to choose members.

OpenStudy (anonymous):

This is the complete question ma'am ..

OpenStudy (loser66):

ok, got it, let me tag someone else @e.mccormick

OpenStudy (anonymous):

thanks :)

OpenStudy (loser66):

I mean: among of 7 members, what is the ratio of Senators and Representatives

OpenStudy (anonymous):

Legislature The Legislative Power resides in the Senate and in the Chamber of Representatives. The Senate consists of 27 members, 2 per electoral district, and 11 elected according to the different districts proportion of population. Two extra seats are granted in each house to the opposition if necessary to limit any party's control to two thirds.

OpenStudy (e.mccormick):

In the house or senate does not matter for the math, so it is a pool of both and picking 7 at random without replacement.

OpenStudy (loser66):

ok, go ahead, friend, I don't understand a word. hihih..

OpenStudy (e.mccormick):

Have you been studying permiutations?

OpenStudy (anonymous):

yes..

OpenStudy (e.mccormick):

Seen the formula for combinations where order does not matter and repetition is not allowed?

OpenStudy (anonymous):

i dont recall..

OpenStudy (e.mccormick):

\[\frac{n!}{r!(n-r)!}\]That spark any memories?

OpenStudy (anonymous):

sure! :)

OpenStudy (e.mccormick):

Got it from there?

OpenStudy (anonymous):

no :(

OpenStudy (e.mccormick):

Well, do you know what the ! means?

OpenStudy (e.mccormick):

Factorials: \[5! = 5\cdot 4\cdot 3\cdot 2\cdot 1\]

OpenStudy (anonymous):

ok got that

OpenStudy (e.mccormick):

If n is the number of items in the pool and your order of selection is not important, but there is not replacement, then the when you choose r items from the pool the possible combinations is given by: \[\frac{n!}{r!(n-r)!}\]

OpenStudy (e.mccormick):

So, what is your total pool size here (n)? ANd how many will you choose from it (r)?

OpenStudy (anonymous):

n=51+27 and r = 7?

OpenStudy (e.mccormick):

Yes. So put those in as 78 and 7, then solve.

OpenStudy (anonymous):

ok thx

OpenStudy (e.mccormick):

So, what did you get?

OpenStudy (anonymous):

78*76*75 and so on?

OpenStudy (e.mccormick):

Well, there is a lot of reduction. The 71 on the bottom nocks out most of it and the 7 nocks out a little more.

OpenStudy (e.mccormick):

\[\frac{78!}{7!(78-7)!}\implies \frac{78!}{7!(71)!}\]See, the 71 on the bottom automatically kills from 71 down to 1. Do you understand what I mean by that?

OpenStudy (e.mccormick):

As an example with smaller numbers: \[\frac{8!}{6!}\implies \frac{8 \cdot 7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1}{6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1}\implies \\ \frac{8 \cdot 7\cdot \not6\cdot \not5\cdot \not4\cdot \not3\cdot \not2\cdot \not1}{\not6\cdot \not5\cdot \not4\cdot \not3\cdot \not2\cdot \not1}\implies \\ 8\cdot 7 = 56\]

OpenStudy (anonymous):

ok i need pen n paper.. i will get back to u tomm. thx

OpenStudy (e.mccormick):

OK... well. I did it in a text file. LOL. But you can see how the 71! makes it a lot simpler. The 7! will help some too. Then you just need to multiply what is left.

OpenStudy (anonymous):

so the answer would be 78times76times75times74times73times72?

OpenStudy (anonymous):

1.3315187e+13 that does seem right

OpenStudy (loser66):

@malia667 honey, I do not follow the steps e.mccormick gave you, but I trust him. you can take it. (sorry for the message I sent you that said "yes , it is" without checking the stuff)

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

Well you are choosing 7 people out of 78. C(78,7). Either use the calculator or apply the formula

OpenStudy (e.mccormick):

Did you remember to also cancel out the 7! part? You are high, and that might do it.

OpenStudy (anonymous):

so what do i multiply?

OpenStudy (e.mccormick):

Because the 71! cancels, you are left simplifying this:\[\frac{78!}{7!71!}\implies \frac{78\cdot 77\cdot 76\cdot 75\cdot 74\cdot 73\cdot 72 }{7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2}\]